OpenStudy (qaisersatti):

i want to ask a simple question for my neice actually i forgot the rules of logrithem solve this question in single expression Log(x-2)log(x+3)log(x+1)-log(x^2-1) please help ...its urgent she has paper tomorrow.

zepdrix (zepdrix):

Like this?? \[\Large \log(x-2)\log(x+3)\log(x+1)-\log(x^2-1)\] Hmm they all have the same base, yes? There should be some type of operator between each log, addition or subtraction, in order for us to simplify this.